Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
mwas - 8:22 Thursday 27 July 2023 (60999) Print this report
Comment to ISC shift: recovery of the ITF + test with RAMS servo ON and OFF (60998)

A few figure to illustrate some of the observations of yesterday.

Figure 1 shows the intereferometer with arms locked and Michelson fringes swinging now, and figure 2 before the RAMS servo isntallation. The top of the fringes seen on B4 is a factor 1.53 higher for the 12MHz and 1.67 higher for the 112MHz. A similar increase is visible with the CITF locked. This contradicts measurment of the modulation depth with the OMC scan, where the 56MHz depth increased from 0.17 to 0.19. So one would expect an increase of only a factor 1.25 for the 112MHz. 

Figure 3, 4 and 5 have the same reference time (purple). Jul 25 at 1:18, before the RAMS servo installation. They all show DRMIT 3F lock data. After the RAMS installation both with (figure 3) and without (figure 4) enabled there is a broad bump between 100Hz and 500Hz on ITF audio channels (in particular B4 and B2). This was not present the last time the RAMS servo was put in place (figure 5).

Do we have an air flow set to high left over somewhere? Lights left on?

The made some single bounce checks to verify that the RAMS servo and new EOM work as expected.

Figure 6 shows in purple OMC locked on the 56MHz USB, with the RAMS loop open, and in red with the RAMS loop closed. The total power on the PD is ~0.8mW in that state. And with the RAMS loop open the fluctuations are at the expected level of a few 1e-7, and with the loop closed the signal is limited by the PD electronic noise. In green is shown with the OMC locked on the 6MHz USB, half of the power on the PD is the tail of the carrier TEM00. The RAMS servo is enabled. In blue is with the OMC stabilized in temperature at the 6MHz frequency and the 6MHz modulation depth reduced by 7dB. Note that moving away the OMC lock from the 6MHz makes this noise that has a pole at ~500Hz shape disappear. So this noise is some interference between the 6MHz and tail of the carrier light. Changing the 8MHz modulation depth by a factor 2 had no impact on it. In any case, it is not much higher than the RAM noise when the servo loop is open, so it shouldn't be a problem for locking.

17:56 (5min) OMC locked on 56MHz USB and B1 PD1/PD2 open. RAMS loop open
18:05 (4min) OMC locked on 56MHz USB and B1 PD1/PD2 open. RAMS loop closed

In between there are seval minutes of data with B1 PD1/PD2 shutter
closed but Vbias enabled.

18:18 (10min) OMC locked on 6MHz USB and B1 PD1/PD2 open. RAMS loop closed

8MHz mod depth increased from 0dBm to 6dBm

18:30:35 (2min) OMC locked on 6MHz USB and B1 PD1/PD2 open. RAMS loop closed. 8\
MHz at 6dBm
 

Figure 7. Note that the change in the spectrum of the power at the output of the PMC have a corresponding vhange in the power reflected by the IMC. And this has been changing by an order of magnitude since at least a week, so is unrelated to the locking troubles.

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×